OpenStudy (anonymous):

Whoever solves these will get MEDAL v+9/3 = 8 2 (n + 5) = -2 -9 x + 1 = -80 -6 = n/2 - 10 - 2 = 2 + v/4 144= - 12 (x + 5)

OpenStudy (anonymous):

@Leader

OpenStudy (anonymous):

@midhun.madhu1987

OpenStudy (anonymous):

@ankit042

OpenStudy (anonymous):

@aussielover101

OpenStudy (anonymous):

Anyone....please help <3

OpenStudy (ankit042):

v= 8-9/3 n+5 =-1 -9x=-81 4=n/2 -4=v/4 -12 =x +5

OpenStudy (ankit042):

I have simplified now you complete :)

OpenStudy (anonymous):

Are they in order from question 1 to question 6

OpenStudy (anonymous):

so now i need to simplify it or it is already good

OpenStudy (anonymous):

5 -6 9 8 -16 -17

OpenStudy (anonymous):

those are the answers? @moralmer000

OpenStudy (anonymous):

yes ma'am

OpenStudy (anonymous):

Thank you so much! :)

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

Is @moralmer000 right about the answers @ankit042

OpenStudy (anonymous):

@ankit042 are u there?

OpenStudy (anonymous):

no hes gone sorry :P